OpenStudy (anonymous):

How to solve 3a+10+4(a+2)-1/2 a by doing the same thing to both sides

OpenStudy (bahrom7893):

is equal to what?

OpenStudy (anonymous):

oh sorry 3a+10+4(a+2) -1/2 a

OpenStudy (anonymous):

3a+10=4(a+2) -1/2 a OK THATS THE RIGHT ONE

OpenStudy (bahrom7893):

okay

OpenStudy (bahrom7893):

3a + 10 = 4(a+2) - 1/2 Divide everything by 4: (3/4) a + 10/4 = a+2 - 1/8

OpenStudy (bahrom7893):

Subtract (3/4) a from both sides: 10/4 = (1/4)a + 2 - 1/8

OpenStudy (bahrom7893):

10/4 = (1/4)a + 15/8 (im just simplifying the right side)

OpenStudy (bahrom7893):

Multiply thru by 4: 10 = a + 15/2 Subtract 15/2 from both sides: 5/2 = a